OpenStudy (anonymous):

Calculate the slope of the tangent to the curve at the given point or value of x. y=1/x , x=1/3

OpenStudy (anonymous):

find the derivative, plug in 1/3

OpenStudy (anonymous):

\[y=\frac{1}{x}\] \[y'=-\frac{1}{x^2}\]

OpenStudy (anonymous):

replace x by 1/3 get \[y'=-\frac{1}{(\frac{1}{3})^2}=-9\]